OpenStudy (anonymous):

How can x2 + 3x + 1 = 2x2 + 2x + 3 be set up as a system of equations?

OpenStudy (anonymous):

y = x2 + 3x + 1 y = 2x2 + 2x + 3 y = 2x2 y = x2 + 2x + 3 y = 2x2 + 2x y = 2x2 + 9 y = x2 y = 2x2 + 2x + 3 which one?